ri2005:

Help! logan found that a slide is 8 feet tall casts a shadow 15 feet long. Then logan notices that a lemon tree cast a shadow 30 feet long how tall is the lemon tree?

faithmiller:

What do you need help with figuring out???

CocoCammron:

How tall is the lemon tree?

ri2005:

yes

CocoCammron:

that is what we are trying to figure out wait one sec

faithmiller:

oh well do you know how he gow 15 for the other one

faithmiller:

I think that it is 23 because i subtracted 7 because he added 7

faithmiller:

what grade r u in

ri2005:

10th

faithmiller:

cool is the answer right

jhonyy9:

Help! logan found that a slide is 8 feet tall casts a shadow 15 feet long. Then logan notices that a lemon tree cast a shadow 30 feet long how tall is the lemon tree? so just use the rule of 3 simple and write in this way 8 feet tall casts a shadow 15 feet long x feet tall casts a shadow 30 feet long cross multiply and calcule the x value x = 30*8/15 = 2*8 = ? ft